# Loss Functions and Metrics in Deep Learning. A Review

### v1

@@ Abstract
One of the essential components of deep learning is the choice of the loss function and performance metrics used to train and evaluate models. This paper reviews the most prevalent loss functions and performance measurements in deep learning. We examine the benefits and limits of each technique and illustrate their application to various deep-learning problems. Our review aims to give a comprehensive picture of the different loss functions and performance indicators used in the most common deep learning tasks and help practitioners choose the best method for their specific task.
@@
Acronyms
AP
Average Precision
AUC-ROC
Area under the Receiver Operating Characteristic curve
BCE
Binary Cross Entropy
BF
Boundary F1 Score
CCE
Categorical Cross Entropy
COCO
Common Objects in Context
FDR
False Discovery Rate
FID
Fréchet Inception Distance
FPR
False Positive Rate
IoU
Intersection Over Union
IS
Inception Score
KL
Kullback-Leibler
MAE
Mean Absolute Error
MAPE
Mean Absolute Percentage Error
MSE
Mean Square Error
NPV
Negative Predictive Value
PQ
Panoptic Quality
PSNR
Peak Signal-to-Noise Ratio
RMSE
Root Mean Square Error
SMAPE
Symmetric Mean Absolute Percentage Error
SSIM
Structural Similarity Index
TDR
True Discovery Rate
TPR
True Positive Rate
VOC
Visual Object Classes
WBCE
Weighted Binary Cross Entropy
YOLO
You Only Look Once
@@ 1 Introduction
Deep Learning has become a powerful tool for solving complex problems in various fields, such as image [1, 2, 3, 4, 5] and speech recognition [6, 7, 8, 9, 10], natural language processing [11, 12, 13, 14, 15], and computer vision [16, 17, 18]. One of the critical components of Deep Learning is the choice of the loss function and performance metrics used to train and evaluate models. Loss functions measure how well a model can approximate the desired output, while performance metrics evaluate the model’s ability to make accurate predictions on unseen data.
Selecting a suitable loss function and performance metric is crucial for achieving good performance in deep learning tasks. However, with a wide range of options available, it can be challenging for practitioners to choose the most appropriate method for their specific task.
This paper aims to comprehensively review the most commonly used loss functions and performance metrics in deep learning. We will discuss the advantages and limitations of each method and provide examples of their application in various Deep Learning tasks. We begin by discussing regression and classification’s most commonly used loss functions, including mean squared error, cross-entropy, and hinge loss. Then, we explain their advantages and limitations and when they are typically used. For example, mean squared error is widely used for regression tasks, while cross-entropy is used for classification tasks. We will also examine more complex tasks such as object detection, segmentation, face recognition, and image generation.
Along the way, we review the most commonly used performance metrics in each category, explaining how these metrics are calculated, their advantages and limitations, and when they are typically used.
@@
A loss function and a performance metric are both used to evaluate the performance of a deep learning model, but they serve different purposes.
A loss function is used during training to optimize the model’s parameters. It measures the difference between the predicted and expected outputs of the model, and the goal of training is to minimize this difference.
On the other hand, a performance metric is used to evaluate the model after training. It measures how well the model can generalize to new data and make accurate predictions. Performance metrics also compare different models or configurations to determine the best-performing one.
The following list describes the common differences between loss functions and performance metrics:
  • Optimization vs. Evaluation: As mentioned previously, loss functions optimize the model’s parameters during training. In contrast, performance metrics evaluate the model’s performance after training.
  • Model-Dependence: Loss functions depend on the model’s architecture and the specific task. Performance metrics, however, are less dependent on the model’s architecture and can be used to compare different models or configurations of a single model.
  • Minimization vs. Maximization: The goal of training a deep learning model is to minimize the loss function. However, evaluating a model aims to maximize the performance metric —except for error performance metrics such as Mean Squared Error.
  • Interpretation: Loss functions can be challenging to interpret as their values are often arbitrary and depend on the specific task and data. On the other hand, performance metrics are often more interpretable as they are used across different tasks.
@@ 2.1 Properties of loss functions
The loss functions have a series of properties that need to be considered when selected for a specific task:
@@
    Convexity: A loss function is convex if any local minimum is the global minimum. Convex loss functions are desirable because they can be easily optimized using gradient-based optimization methods.
@@
    Smoothness: Loss function should have a continuous gradient and no sharp transitions or spikes.
@@
    Multi-modality: A loss function is considered multi-modal if it has multiple global minima. Multi-modal loss functions can be useful for tasks requiring the model to learn multiple data representations.
  7. 7.
    Monotonicity: A loss function is monotonous if its value decreases as the predicted output approaches the true output. Monotonicity ensures that the optimization process is moving toward the correct solution.
  8. 8.
    Invariance: A loss function is invariant if it remains unchanged under particular input or output transformations. Invariance is valuable when working with data that may be transformed in various ways, such as rotation, scaling, or translation.
The following sections review the loss functions and performance metrics for common deep learning tasks. Table 1 summarizes common vision-related tasks with their common loss functions used and performance metrics.
Table 1:
Loss functions and performance metrics for deep learning tasks.
Deep Learning Task
Loss Functions
Performance Metrics
Regression
MSE (3.1.1)
@@
Multi-Class Classification
CCE (4.1.3)
@@
Object Detection
Smooth L1 (5.1.1)
@@
Face Recognition
A-Softmax (7.1.2)
@@
Image Generation
Adversarial Loss (8.1.3)
@@
3 Regression
Regression is a supervised learning problem in machine learning that aims to predict a continuous output value based on one or more input features. Common regression models include linear regression, polynomial regression, and regression trees.
Linear regression assumes a linear relationship between the independent and dependent variables. It is represented by the equation
$\hat{y}=\beta_{0}+\beta_{1}x_{1}+\beta_{2}x_{2}+\dots+\beta_{n}x_{n},$ (1)
where $\hat{y}$ is the predicted value, $\beta_{0}$ is the intercept or bias, $\beta_{1},\beta_{2},...,\beta_{n}$ are the coefficients or weights corresponding to the input features or independent variables $x_{1},x_{2},...,x_{n}$ and $n$ is the number of input features.
The goal is to find the bias and the coefficient values that minimize the difference between the predicted and actual values, usually using a loss function such as Mean Squared Error (MSE) or Mean Absolute Error (MAE).
In polynomial regression, the relationship between the independent variable $x$ and the dependent variable $y$ is modeled as an $n^{th}$ degree polynomial. This is useful for capturing complex, non-linear relationships between input and output variables. The general form of a polynomial regression equation is given by
$\hat{y}=\beta_{0}+\beta_{1}x+\beta_{2}x^{2}+\beta_{3}x^{3}+\dots+\beta_{n}x^{n},$ (2)
where $\hat{y}$ is the predicted value, $\beta_{0}$ is the intercept or bias, $\beta_{1},\beta_{2},...,\beta_{n}$ are the coefficients corresponding to the powers of $x$ and $n$ is the degree of the polynomial.
Like linear regression, the objective is to find the bias and the coefficients that minimize the difference between the predicted and the actual values. However, high-degree polynomials tend to overfit when the model becomes excessively complex such that it performs well on training data but poorly on unseen or test data.
Regression trees, on the other hand, are a type of decision tree where the output is a continuous variable. Unlike linear and polynomial regression models that establish a single prediction equation, regression trees split the input space into smaller regions where a simple model is used. The tree is built during training through a process known as binary recursive partitioning. The output for a new instance is predicted by traversing the tree until a leaf node is reached. The value associated with the leaf node is typically the mean target value of the training samples in this node. Unlike polynomial regression, this model can capture complex, non-linear relationships and interactions between features without specifying them explicitly. However, regression trees can also overfit the training data if not properly pruned or controlled, leading to poor generalization performance on new, unseen data. Figure 1 shows a regression tree.
Figure 1: Regression Tree.
Regression is used in various domains, including finance, healthcare, social sciences, sports, and engineering. Some practical applications include house price prediction [19], energy consumption forecasting [20], healthcare and disease prediction [21], stock price forecasting [22], and customer lifetime value prediction [23].
In the following subsections, we will review the most common lost functions and performance metrics used for regression.
3.1 Regression Loss Functions
@@
Loss Function
Applications
@@
Linear Regression, Ridge Regression, Lasso Regression, Neural Networks, Support Vector Regression, Decision Trees, Random Forests, Gradient Boosting
Mean Absolute Error (MAE)
Quantile Regression, Robust Regression, L1 Regression, Neural Networks, Decision Trees, Random Forests, Gradient Boosting
Huber Loss
Robust Linear Regression, Robust Neural Networks, Gradient Boosting, Random Forests
Log-Cosh Loss
Robust Regression, Neural Networks, Gradient Boosting
Quantile Loss
Quantile Regression, Distributional Regression, Extreme Value Prediction
Poisson Loss
Poisson Regression, Count Data Prediction, Generalized Linear Models, Neural Networks, Gradient Boosting
3.1.1 Mean Squared Error (MSE)
The Mean Square Error (MSE) measures the average of the squared differences between the predicted values and the true values [24]. The MSE loss function can be defined mathematically as
@@ 3.1.1 Mean Squared Error (MSE)
  • Non-negative: Since the differences between the predicted and actual values are squared, MSE is always non-negative. A value of 0 indicates a perfect fit, while larger values correspond to higher discrepancies between predictions and actual values.
  • Quadratic: MSE is a quadratic function of the prediction errors, which means it places more emphasis on larger errors than smaller ones. This property makes it sensitive to outliers and can lead to models that prioritize reducing large errors over smaller ones.
  • Differentiable: MSE is a smooth and continuous function for the model parameters. This property allows for the efficient computation of gradients, which is essential for optimization algorithms like gradient descent.
@@
  • Scale-dependent: The value of MSE depends on the scale of the target variable, making it difficult to compare the performance of models across different problems or target variable scales. For this purpose, researchers often use the root mean squared error (RMSE) or mean squared percentage error (MSPE).
The MSE, also called L2 loss, is computationally simple. However, it is not robust to outliers due to the square of the error term. Thus if the data includes outliers, it is better to use another loss function, such as Mean Absolute Error (MAE) which is more robust to outliers, or Huber Loss, which is a combination of MSE and MAE. The MSE is also used as a performance metric.
3.1.2 Mean Absolute Error (MAE)
The Mean Absolute Error (MAE) is another commonly used loss function in regression problems. It measures the average of the absolute differences between the predicted values and the true values [25]. The MAE loss can be defined as
@@ 3.1.2 Mean Absolute Error (MAE)
  • Linear: MAE is a linear function of the prediction errors, which treats all errors equally regardless of their magnitude. This property makes MAE less sensitive to outliers than MSE, as it does not disproportionately emphasize large errors.
  • Robust: Due to its linear nature and reduced sensitivity to outliers, MAE is considered a more robust loss function than MSE. This makes it suitable for applications where the presence of outliers is expected or the distribution of errors is not symmetric.
@@
  • Scale-dependent: Like MSE, the value of MAE depends on the scale of the target variable, making it difficult to compare the performance of models across different problems or target variable scales. To address this issue, researchers often use scale-invariant metrics such as mean absolute percentage error (MAPE) or normalized mean absolute error (NMAE) to compare models across different scales or units.
The MAE, called L1 loss, is often used as an evaluation metric. It is computationally simple and easy to understand, but it does not have the smooth and differentiable property of the MSE and is not sensitive to outliers.
3.1.3 Huber Loss
@@ 3.1.4 Log-Cosh Loss
3.1.5 Quantile Loss
@@
$L(y,\hat{y})=q\cdot\max(y-\hat{y},0)+(1-q)\cdot\max(\hat{y}-y,0)$ (7)
$\max(a,b)$ represents the maximum of $a$ and $b$. The expression $y-\hat{y}$ is used when the prediction underestimates, and $\hat{y}-y$ is used when the prediction overestimates. The loss is scaled by $q$ for underestimations and $(1-q)$ for overestimations.
@@ 3.1.6 Poisson Loss
3.2 Regression Performance Metrics
@@
Performance Metric
Applications
Mean Squared Error (MSE)
General-purpose regression, model selection,
@@
Root Mean Squared Error (RMSE)
General-purpose regression, model selection,
@@
Mean Absolute Error (MAE)
General-purpose regression, model selection,
@@
R-squared (R²)
Model evaluation, goodness-of-fit, linear regression,
@@
Adjusted R-squared
Model evaluation, goodness-of-fit, linear regression,
@@
Mean Squared Logarithmic Error
Forecasting, model evaluation, skewed
target distributions, finance, sales prediction
Mean Absolute Percentage Error
Forecasting, model evaluation, time series analysis,
business analytics, supply chain optimization
3.2.1 Root Mean Squared Error (RMSE)
The Root Mean Square Error (RMSE) is the square root of the mean squared error (MSE) defined as
@@ 3.2.1 Root Mean Squared Error (RMSE)
3.2.2 Mean Absolute Percentage Error (MAPE)
The Mean Absolute Percentage Error (MAPE) measures the average percentage error of the model’s predictions compared to the true values. It is defined as
@@ 3.2.2 Mean Absolute Percentage Error (MAPE)
One of the advantages of using MAPE is that it is easy to interpret, as it is expressed in percentage terms. It is also scale-independent, which can be used to compare models across different scales of the target variable. However, it has two limitations: it can produce undefined results when $y_{i}$ is zero and is sensitive to outliers.
3.2.3 Symmetric Mean Absolute Percentage Error (SMAPE)
The Symmetric Mean Absolute Percentage Error (SMAPE) is a variation of the Mean Absolute Percentage Error (MAPE) commonly used to evaluate the accuracy of predictions in time series forecasting [49]. SMAPE is defined as
@@ 3.2.3 Symmetric Mean Absolute Percentage Error (SMAPE)
3.2.4 Coefficient of Determination $R^{2}$
The Coefficient of Determination ($R^{2}$), measures how well the model can explain the variation in the target variable [55]. $R^{2}$ is defined as the proportion of the variance in the target variable that the model explains. It ranges from 0 to 1, where 0 means that the model does not explain any variation in the target variable, and one means that the model explains all the variation in the target variable.
The formula for R-squared is
$R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}i)^{2}}{\sum{i=1}^{n}(y_{i}-\bar{y})^{2}},$ (14)
@@ 3.2.4 Coefficient of Determination $R^{2}$
    Measures the relationship between the model and the response variable: R-squared describes the strength of the relationship between the model and the response variable on a convenient 0 – 1 scale.
@@
    Interpretable: It can be more interpretable than other statistics because it provides a percentage that can be intuitively understood.
@@
    Helps in model selection: If we have two models, we can compare their R-squared values as a part of the selection process. The model with the higher R-squared could indicate a better fit.
@@
    Misleading with non-linear relationships: $R^{2}$ works as intended in a simple linear regression model with one explanatory variable but can be misleading with more complex, nonlinear, or multiple regression models.
@@
    Influenced by the number of predictors: $R^{2}$ always increases as we add more predictors to a model, even if they are unrelated to the outcome variable. This can lead to overly complex models that overfit the data. This is the benefit of the adjusted $R^{2}$, which adjusts the $R^{2}$ value based on the number of predictors in the model.
@@
    Sensitive to outliers: $R^{2}$ is sensitive to outliers.
@@
    Does not check for biased predictions: $R^{2}$ cannot determine whether the coefficient estimates and predictions are biased, which is to say, whether the predictions systematically over or underestimate the actual values.
@@
    Limitation with small sample sizes: When the sample size is small, the $R^{2}$ value might be unreliable. It can be artificially high or low and might not represent the true strength of the relationship between the variables.
3.2.5 Adjusted $R^{2}$
Adjusted $R^{2}$ is a modified version of $R^{2}$ that has been adjusted for the number of predictors in the model. It increases only if the new term improves the model more than would be expected by chance. It decreases when a predictor improves the model by less than expected by chance [56]. The adjusted R-squared is defined as
@@ 3.2.5 Adjusted $R^{2}$
where $n$ is the number of observations, $k$ is the number of predictors. The adjustment is a penalty for adding unnecessary predictors to the model. This penalty increases with the increase in the number of predictors. This is particularly useful in multiple regression, where several predictors are used simultaneously.
The Adjusted $R^{2}$ is often used for model comparison, as it won’t necessarily increase with adding more variables to the model, unlike regular $R^{2}$. It is useful when we need to compare models of different sizes. Unlike $R^{2}$, its value can be negative, meaning that the model is a poor fit for the data.
4 Classification

### v3

@@ Abstract
When training or evaluating deep learning models, two essential parts are picking the proper loss function and deciding on performance metrics. In this paper, we provide a comprehensive overview of the most common loss functions and metrics used across many different types of deep learning tasks, from general tasks such as regression and classification to more specific tasks in Computer Vision and Natural Language Processing. We introduce the formula for each loss and metric, discuss their strengths and limitations, and describe how these methods can be applied to various problems within deep learning. We hope this work serves as a reference for researchers and practitioners in the field, helping them make informed decisions when selecting the most appropriate loss function and performance metrics for their deep learning projects.
@@ 1 Introduction
Deep Learning has become the dominant technology for solving problems involving unstructured data, such as images. [1, 2, 3, 4, 5, 6, 7, 8], video, audio [9, 10, 11, 12, 13], and text [14, 15, 16, 17, 18]. One of the critical components of deep learning is the selection of the loss function and performance metrics used for training and evaluating models. Loss functions measure how effectively a model can approximate the desired output, while performance metrics assess the model’s ability to make accurate predictions on unseen data. Choosing the appropriate loss function and performance metric is essential for success in deep learning tasks. However, with many options to choose from, it can be challenging for practitioners to determine the most suitable method for their specific task.
In this paper, we present a thorough overview of the most commonly utilized loss functions and performance metrics in the field of deep learning. We analyze the strengths and weaknesses of each method and provide illustrative examples of their application across various deep learning tasks.
First, we delve into the prevalent regression and classification loss functions, such as mean squared error, cross-entropy, and hinge loss, delineating their respective advantages, limitations, and typical use cases. Subsequently, we explore standard tasks in computer vision, such as image classification, object detection, image segmentation, face recognition, and image generation. Finally, we conclude our review by outlining the prevalent loss functions and metrics employed in natural language processing.
This paper is structured as follows. Section 2 outlines the difference between loss functions and performance metrics. Section 3 provides an overview of the most widely used losses and metrics in regression and classification. Section 4 delves into the prevalent computer vision tasks, detailing the associated loss functions and metrics. Section 5 concentrates on the use of losses and metrics in NLP. Section 6 presents a discussion of this research and, finally, Section 7 rounds out the paper with a concluding statement.
@@
Loss functions and performance metrics are two distinct tools for evaluating the performance of a deep learning model and serve different purposes.
During training, a loss function is used to optimize the model’s parameters. Measures the difference between the predicted and expected outputs of the model. The objective of training is to minimize this difference.
In contrast, a performance metric is used to evaluate the model after training. It helps to determine how well the model can generalize to new data and make accurate predictions. Performance metrics also aid in comparing different models or configurations to identify the best-performing one.
The following list details the key differences between loss functions and performance metrics:
  • During the training of a deep learning model, loss functions are used to optimize the model’s parameters, whereas performance metrics are used to evaluate the model’s performance after training.
  • The choice of loss function typically depends on the model’s architecture and the specific task at hand. In contrast, performance metrics are less dependent on the model’s architecture and can be used to compare different models or configurations of a single model.
  • The ultimate goal of training a deep learning model is to minimize the loss function, while evaluating a model aims to maximize the performance metric, with the exception of error performance metrics such as Mean Squared Error.
  • Loss functions can be challenging to interpret as their values are often arbitrary and depend on the specific task and data. In contrast, performance metrics are often more interpretable and can be used across different tasks.
@@ 2.1 Properties of loss functions
Loss functions have a series of properties that need to be considered when selected for a specific task:
@@
    Convexity: A loss function is convex if any local minimum is also the global minimum. Convex loss functions are desirable because they can be easily optimized using gradient-based optimization methods.
@@
    Smoothness: A loss function should have a continuous gradient and no sharp transitions or spikes.
@@
    Monotonicity: A loss function is monotonic if its value decreases as the predicted output approaches the true output. Monotonicity ensures that the optimization process is moving toward the correct solution.
The tables below provide a summary of the loss functions and performance metrics that are discussed in this work. Table 1 outlines the functions and metrics employed in general tasks such as regression, binary classification, and multiclass classification. Table 2 provides an overview of the loss functions and metrics related to computer vision, and table 3 summarizes the work on natural language processing.
Table 1: Loss functions and performance metrics for general tasks.
Deep Learning Task Loss Functions Performance Metrics
Regression MSE (3.1.1)
@@
Multi-Class Classification CCE (3.2.1)
@@
Table 2: Loss functions and performance metrics used in Computer Vision.
Deep Learning Task Loss Functions Performance Metrics
Object Detection Smooth L1 (4.2.1)
@@
Face Recognition A-Softmax (4.4.1)
@@
Image Generation Adversarial Loss (4.5.1)
@@
Table 3: Loss functions and performance metrics used in Natural Language Processing.
Deep Learning Task Loss Functions Performance Metrics
Text Classification CE (3.2.1)
Hinge loss (5.1.2)
Accuracy (5.3.1)
P/R/F1 (5.3.2)
AUC-ROC (5.3.3)
Language Modeling T-CE (5.1.1) Perplexity (5.3.6)
BLEU (5.3.4)
ROUGE (5.3.5)
Machine Translation T-CE (5.1.1)
MRT (5.2.1)
REINFORCE (5.2.2)
BLEU (5.3.4)
ROUGE (5.3.5)
Perplexity (5.3.6)
Exact match (5.3.7)
Name Entity Recognition T-CE (5.1.1) Accuracy (5.3.1)
P/R/F1 (5.3.2)
Exact match (5.3.7)
Part-of-Speech Tagging T-CE (5.1.1) Accuracy (5.3.1)
P/R/F1 (5.3.2)
Sentiment Analysis CE (3.2.1)
Hinge loss (5.1.2)
Cosine Similarity (5.1.3)
Adj. Circle (4.4.1)
P/R/F1 (5.3.2)
AUC-ROC (5.3.3)
Text Summarization T-CE (5.1.1)
MRT (5.2.1)
REINFORCE (5.2.2)
BLEU (5.3.4)
ROUGE (5.3.5)
Exact match (5.3.7)
Question Answering CE (3.2.1)
Hinge loss (5.1.2)
Cosine Similarity
P/R/F1 (5.3.2)
Exact match (5.3.7)
Language Detection CE (3.2.1)
Hinge loss (5.1.2)
Accuracy (5.3.1)
P/R/F1 (5.3.2)
3 Basic Tasks
In the following sections, we dive into two essential tasks in machine learning: regression and classification. We examine the loss functions and performance metrics used and offer practical guidance on when to use each, as well as real-world applications.
3.1 Regression
Regression is a supervised learning problem in machine learning that aims to predict a continuous output value based on one or more input features. Regression is used in various domains, including finance, healthcare, social sciences, sports, and engineering. Some practical applications include house price prediction [19], energy consumption forecasting [20], healthcare and disease prediction [21], stock price forecasting [22], and customer lifetime value prediction [23].
In the following subsections, we review the most common lost functions and performance metrics used for regression.
3.1.1 Regression Loss Functions
@@
Loss Function Applications
Mean Squared Error (MSE) Linear Regression, Ridge Regression, Lasso Regression, Neural Networks, Support Vector Regression, Decision Trees, Random Forests, Gradient Boosting
Mean Absolute Error (MAE) Quantile Regression, Robust Regression, L1 Regression, Neural Networks, Decision Trees, Random Forests, Gradient Boosting
Huber Loss Robust Linear Regression, Robust Neural Networks, Gradient Boosting, Random Forests
Log-Cosh Loss Robust Regression, Neural Networks, Gradient Boosting
Quantile Loss Quantile Regression, Distributional Regression, Extreme Value Prediction
Poisson Loss Poisson Regression, Count Data Prediction, Generalized Linear Models, Neural Networks, Gradient Boosting
@@
The MSE measures the average of the squared differences between the predicted values and the true values [24]. The MSE loss function can be defined mathematically as
@@
  • Non-negative: MSE is always non-negative because the differences between the predicted and actual values are squared. A value of 0 indicates a perfect fit, while larger values correspond to higher discrepancies between predictions and actual values.
  • Sensitive to outliers: MSE is a quadratic function of the prediction errors, which means it places more emphasis on larger errors than smaller ones. This property makes it sensitive to outliers and can lead to models that prioritize reducing large errors over smaller ones.
  • Differentiable: MSE is a smooth and continuous function for the model parameters. This property allows for the efficient computation of gradients, which is essential for optimization algorithms such as gradient descent.
@@
  • Scale-dependent: The value of MSE depends on the scale of the target variable, making it difficult to compare the performance of models across different problems or target variable scales. For this purpose, researchers often use the root mean squared error (RMSE) or mean squared percentage error (MSPE) for computing the performance of the model.
The Mean Squared Error (MSE), which is also known as L2 loss, is a simple and widely used method for calculating the average of the squared differences between predicted and actual values. It is not only utilized as a loss function during the training of models but also serves as a performance metric to evaluate the accuracy of predictions. However, it is important to note that the use of the square of the error term can make it susceptible to outliers in the data. In such cases, it is recommended to use other loss functions such as Mean Absolute Error (MAE) or Huber Loss, which are more robust to outliers.
Mean Absolute Error (MAE)
The MAE is another commonly used loss function in regression problems. It measures the average of the absolute differences between the predicted values and the true values [25]. The MAE loss can be defined as
@@
  • Robust to outliers: MAE is a linear function of the prediction errors, which treats all errors equally regardless of their magnitude. This property makes MAE less sensitive to outliers than MSE, as it does not disproportionately emphasize large errors. MAE is considered a more robust loss function than MSE. This makes it suitable for applications where the presence of outliers is expected or the distribution of errors is not symmetric.
@@
The MAE called L1 loss, is often used as an evaluation metric. It is computationally simple and easy to understand, but it does not have the smooth and differentiable property of the MSE and is not sensitive to outliers. However, like MSE, the value of MAE depends on the scale of the target variable, making it difficult to compare the performance of models across different problems or target variable scales. To address this issue, it is often use scale-invariant metrics such as mean absolute percentage error (MAPE) or normalized mean absolute error (NMAE) to compare models across different scales or units.
Huber Loss
@@
Huber Loss vs Log-Cosh Loss
Huber loss and Log-Cosh loss are both loss functions that aim to improve robustness to outliers compared to the standard Mean Squared Error (MSE). They do so by blending properties of MSE and Mean Absolute Error (MAE), but each has distinct characteristics and advantages.
We use Huber loss when we need robustness to outliers and prefer a loss function that has a controllable transition point (via the $\delta$ parameter) between quadratic and linear behavior. It is useful when we have a reason to define a specific point where the loss function should switch from quadratic to linear, depending on the noise characteristics of the data.
We use Log-Cosh Loss when we want a smooth and more universally applicable loss function that is always differentiable, making it suitable for optimization algorithms that benefit from smooth gradients. It can be a better choice when we do not have clear reasons to manually set a transition threshold as in Huber loss.
Quantile Loss
@@
$L(y,\hat{y})=q\cdot\max(y-\hat{y},0)+(1-q)\cdot\max(\hat{y}-y,0),$ (5)
where $\max(a,b)$ represents the maximum of $a$ and $b$. The expression $y-\hat{y}$ is used when the prediction underestimates, and $\hat{y}-y$ is used when the prediction overestimates. The loss is scaled by $q$ for underestimations and $(1-q)$ for overestimations.
@@
3.1.2 Regression Performance Metrics
@@
Performance Metric Applications
Mean Squared Error (MSE) General-purpose regression, model selection,
@@
Root Mean Squared Error (RMSE) General-purpose regression, model selection,
@@
Mean Absolute Error (MAE) General-purpose regression, model selection,
@@
R-squared (R²) Model evaluation, goodness-of-fit, linear regression,
@@
Adjusted R-squared Model evaluation, goodness-of-fit, linear regression,
@@
Mean Squared Logarithmic Error Forecasting, model evaluation, skewed
(MSLE) target distributions, finance, sales prediction
Mean Absolute Percentage Error Forecasting, model evaluation, time series analysis,
(MAPE) business analytics, supply chain optimization
Root Mean Squared Error (RMSE)
The RMSE is the square root of the mean squared error (MSE) defined as
@@
The RMSE can be interpreted geometrically as the Euclidean distance between the vector of observed values and the vector of predictions, which is analogous to the length of a hypotenuse in a right triangle, representing the disparity between predicted and actual values.
RMSE is widely used in fields where accurate error magnitude is critical, such as in financial forecasting, energy load predictions, and more generally in regression analysis and machine learning. This is due to its ability to penalize large errors more severely, which can be crucial for some applications.
The limitations of the RMSE are two-fold: on the one hand, it is sensitive to outliers; on the other hand, RMSE does not distinguish between types of errors (systematic and random). The scale and size of the dataset also influences it:
[…]
where $\sigma_{y}$ is the standard deviation of the observed values.
To make RMSE comparable across different datasets or models, it can be normalized:
Normalized $\text{Normalized RMSE}=\frac{RMSE}{\sigma_{y}},$ (11)
which standardizes the RMSE by the variability of the dataset, thereby removing scale effects.
Interpreting RMSE involves comparing it to the standard deviation of the dependent variable. A lower RMSE relative to the standard deviation suggests a model with a better fit:
Relative $\text{Relative Error}=\frac{RMSE}{\sigma_{y}},$ (12)
which helps assess model accuracy relative to the inherent variability of the data.
Mean Absolute Percentage Error (MAPE)
The MAPE measures the average percentage error of the model’s predictions compared to the true values. It is defined as
@@
One of the advantages of using MAPE is that it is easy to interpret, as it is expressed in percentage terms. It is also scale-independent, which allows comparison of forecast accuracy across different scales of the target variable. However, MAPE has two main limitations:
  • It can produce undefined results when $y_{i}=0$, which may require modifying the dataset or the formula to handle such cases.
  • It is sensitive to outliers because large errors in the denominator can disproportionately influence the overall percentage error.
MAPE is particularly useful in scenarios where the relative error is more important than the absolute error magnitude. It is extensively used in industries like finance and retail where percentage errors provide a more intuitive sense of prediction errors relative to scales. Moreover, it is beneficial in comparative model assessments across datasets where absolute scales of targets vary widely.
Symmetric Mean Absolute Percentage Error (SMAPE)
The SMAPE is a variation of the Mean Absolute Percentage Error (MAPE) commonly used to evaluate the accuracy of predictions in time series forecasting [49]. SMAPE is defined as
@@
Coefficient of Determination $R^{2}$
The Coefficient of Determination ($R^{2}$), measures how well the model explains the variation in the target variable [55]. $R^{2}$ is defined as the proportion of the variance in the target variable that the model explains, ranging from 0 to 1. A value of 0 indicates that the model does not explain any variation in the target variable, while a value of 1 means it explains all the variation. It is possible for $R^{2}$ to be negative, which indicates that the model performs worse than a simple mean of the target variable.
The formula for R-squared is:
$R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}},$ (15)
@@
    Measures the relationship between the model and the response variable: It describes the strength of the relationship on a scale from 0 to 1, which is intuitively understandable.
@@
    Provides a percentage indicating how much of the variability in the response variable can be explained by the model.
@@
    Allows comparison of models; higher $R^{2}$ values generally indicate a better fit, though this should be used cautiously.
@@
    It is best suited for linear models, can be misleading for non-linear models or complex relationships.
@@
    Can increase with the number of predictors, potentially leading to overfitting. Adjusted $R^{2}$ is often preferred in multiple regression contexts to account for this.
@@
    Extreme values can significantly impact $R^{2}$.
@@
    Can not tell if predictions are systematically over or underestimate actual values.
@@
    May not accurately reflect the strength of a relationship in small datasets.
Adjusted $R^{2}$
Adjusted $R^{2}$ is a modified version of $R^{2}$ that has been adjusted for the number of predictors in the model. It increases only if the new term improves the model more than would be expected by chance, and it decreases when a predictor improves the model by less than expected by chance [56]. The adjusted $R^{2}$ is defined as
@@
where $n$ is the number of observations and $k$ is the number of predictors. This adjustment serves as a penalty for adding unnecessary predictors to the model, with the penalty increasing as the number of predictors increases. Consequently, adjusted $R^{2}$ provides a more accurate assessment of model fit, particularly in multiple regression analyses where several predictors are used simultaneously.
Adjusted $R^{2}$ is commonly used for model comparison because it does not necessarily increase with the addition of more variables, unlike regular $R^{2}$. It allows for fair comparisons between models of different sizes by penalizing overfitting and rewarding models that achieve a better fit without unnecessarily increasing complexity.
Interpreting adjusted $R^{2}$ values is similar to interpreting regular $R^{2}$, with positive values indicating the proportion of variance explained by the model after adjusting for the number of predictors. However, negative values indicate that the model is worse than a model that simply predicts the mean of the dependent variable.
For example, consider a regression model predicting housing prices using various predictors such as square footage, number of bedrooms, and location. The adjusted $R^{2}$ value would help determine whether adding additional predictors improves the model’s predictive power or simply adds unnecessary complexity.
3.2 Classification